THE DISTANCE BETWEEN STATION ATENA AND STATION BARCENA IS 84 MILES. A TRAIN STARTS FROM ATENA TOWARDS BARCENA. A BIRD STARTS AT THE SAME TIME FROM BARCENA STRAIGHT TOWARDS THE MOVING TRAIN. ON REACHING THE TRAIN, IT INSTANEOUSLY TURNS BACK AND RETURNS BARCENA. THE BIRD MAKES THESE JOURNEYS FROM BARCENA TO THE TRAIN AND BACK TO BARCENA CONTINUOUSLY TILL THE TRAIN REACH BARCENA. THE BIRD FINALLY RETURNS TO BARCENA AND REST. CALCULATE THE TOTAL DISTANCE IN MILES THE BIRD TRAVELS IN THE FOLLOWING CASE:
THE BIRD FLIES AT 80 MILES PER HOUR AND THE SPEED OF THE TRAIN IS 60 MILES PER HOUR.

Answer by mananth(16949)

speed of train = 60 mph
distance it travels to Barcena = 84 miles
time it travels = 84/60 = 14/10 hours.
..
Irrespective of where it flies the bird keeps flying only for 14/10 hours.
That is the time the train takes to reach barcena.
distance the bird travels = speed * time
speed of bird = 80mph
= 80*14/10 hours
= 8*14
=112 miles
the bird flies for 112 miles